### v2

# Towards Active Synthetic Data Generation for Finetuning Language Models

Abstract
A common and effective means for improving language model capabilities involves finetuning a “student” language model’s parameters on generations from a more proficient “teacher” model. Termed “synthetic data”, these generations are often produced before any student finetuning, but some work has considered generating new synthetic samples as training progresses. This paper studies and advocates for the latter case, where data are generated in an iterative, closed-loop fashion that is guided by the current state of the student model. For a fixed budget of generated samples, or a budget in terms of compute spent querying a teacher, we show that this curation of finetuning data affords improved student performance over static generation. Further, while there have been several LLM-specific methods proposed that operate in this regime, we find that simple, inexpensive selection criteria from the active learning literature tend to be most performant. We validate these claims across four mathematical and logical reasoning datasets using four different small language models.
1 Introduction
Despite the tremendous cost of inference, Large Language Models (LLMs) have risen to prominence as a result of their remarkable abilities across a wide array of reasoning and factual knowledge tasks (Achiam et al., 2023; Bubeck et al., 2023; Katz et al., 2024). As agentic systems capable of interacting with the external world emerge, these models are poised to become even more ubiquitous in science, technology, and society, but the tremendous inference cost presents a challenge for realizing the full potential of these agents.
One way to quell the computational expense associated with LLM inference is to use small language models (SLMs). With orders of magnitude fewer parameters, SLMs are faster, cheaper, and easier to finetune for specialised skills like tool use, making them natural specialists using proprietary data or within bespoke agentic systems (Belcak et al., 2025).
Training language models typically involves three stages: pre-training on large, general-purpose corpora, supervised finetuning (SFT), and reinforcement learning from human feedback (RLHF) or from verifiable rewards (RLVR) (Ouyang et al., 2022). SFT, the focus of this work, is critical for
Figure 1: Overview of iterative synthetic data generation (Algorithm 1). The student model guides synthetic data generation by prioritizing which data are used as an example for the teacher model to generate a new synthetic data point (Section 4.2). The student finetunes on teacher generated synthetic instruction data.
adapting a base model to a target distribution, and is especially common when training SLMs to improve their task-specific performance.
In practice, real-world data for SFT can be hard to obtain, or may lack desirable properties such as chain-of-thought reasoning (Wei et al., 2022). Consequently, a typical strategy involves synthesizing a corpus of question and answer pairs from a larger, more capable model (Mitra et al., 2024; Liu et al., 2024a). This process usually begins with a small seed dataset, which a teacher LLM uses to produce supplementary synthetic samples before the student SLM is finetuned on the resulting sequences.
Still, evidence suggests that generating a large, static synthetic dataset is frequently wasteful, as it can often be drastically pruned with little to no degradation in trained model capabilities  (Chen et al., 2023; Zhou et al., 2024). As such, this paper explores an iterative, targeted approach to synthetic data generation that is student-aware and improves data efficiency—achieving stronger performance under a fixed data generation budget than naive, static generation—thereby yielding a superior performance–training-set-size Pareto frontier (see Section 2 for a formal definition).
To facilitate productive learning, this work studies how we can effectively cater to the state of the student model and guide synthetic data generation by a teacher LLM via prompting (Mitra et al., 2024; Liu et al., 2024a; Luo et al., 2025). This results in an iterative scheme, where the updated student can be reused to guide further teacher-generated samples (Figure 1). Prior work has considered this paradigm by prioritizing incorrect student answers (Lee et al., 2024) and using LLM-as-a-judge scoring (Jiang et al., 2023b), but they do not draw upon the vast active learning and data selection literature. Instead, this paper advocates for the generation of data that are conditioned on samples that have been prioritized by an active learning algorithm. The resulting dataset enables more effective and data efficient finetuning of the SLM student model (see Section 5.4 for evidence supporting this claim).
Our work makes the following contributions:
  • We provide a benchmark study for iterative synthetic data generation rooted in prior work on active learning and data selection, and compare to static synthetic dataset generation. We show improvements in data efficiency when comparing to generating a single large synthetic instruction dataset, which is a typical approach to student post-training (Mitra et al., 2024; Luo et al., 2025).
  • We compare a range of methods for selecting samples for seeding synthetic data generation, including those that favour uncertainty, diversity, or difficulty. We conclude that simple methods rooted in active learning, such as using the loss of the student’s own prediction are most data efficient. In contrast, expensive and contemporary methods that use an LLM to judge the difficulty and quality of data, i.e. LLM-as-a-judge (Zheng et al., 2023; Jiang et al., 2023b), surprisingly underperform when compared to simpler alternatives.
  • We show that synthetic data generation is, to a certain extent, steerable: properties of teacher-generated synthetic data resemble those used to seed the generation process. If the student selects challenging data—measured as samples that induce high student loss—the teacher generates data with correspondingly high loss on average. Given this relationship, and the fact that recent, specialised, LLM-based strategies often fall short, we argue that research in selection strategies is a fruitful and underexplored avenue for advancing the performance of small language models.
2 Preliminaries
Notation.
[…]
$\mathcal{L}(z,{\bm{\theta}})=\nicefrac{{-1}}{{|y|}}\sum_{j=1}^{|y|}y_{j}\log f_{{\bm{\theta}}}(x,y_{<j}).$
The model $f_{{\bm{\theta}}}(\cdot)$ autoregressively generates the next token $\hat{y}_{j}=f_{{\bm{\theta}}}(x,\hat{y}_{<j})$ in the sequence.
Data Efficiency.
For a fixed number of samples, if better generalization performance can be achieved by training on one subset of a larger dataset than on another, the former can be considered more data efficient. Formally, let $P$ be the true data distribution over our data $z=(x,y)$. For a selection algorithm $\phi$ that produces a dataset $S_{n}^{\phi}=\left\{z_{i}\right\}_{i=1}^{n}\overset{\phi}{\sim}P$, model parameters ${\bm{\theta}}_{n}^{\phi}$ result from minimizing the loss over $S_{n}^{\phi}$. We define the performance, accuracy for example, on a single sample as
$\text{perf}_{\phi}(z,{\bm{\theta}}^{\phi}_{n})=\bm{1}\left\{y=f_{{\bm{\theta}}_{n}^{\phi}}(x)\right\},$ (1)
and the expected performance as
$\text{perf}_{\phi}(n)=\mathbb{E}_{z\sim P}\mathbb{E}_{S_{n}^{\phi}\sim P}\left[\text{perf}_{\phi}(z,{\bm{\theta}}_{n}(S_{n}^{\phi}))\right].$ (2)
Assuming a monotonic increase in performance with $n$, for some target performance $\tau$, the sample complexity can be defined as
$N_{\phi}(\tau)=\inf\left\{n:\text{perf}_{\phi}(n)\geq\tau\right\},$ (3)
which measures the smallest $n$ such that $\text{perf}_{\phi}(n)\geq\tau$. For a fixed architecture $f(\cdot)$, algorithm $\alpha$ is more data-efficient than algorithm $\beta$ at level $\tau$ only if $N_{\alpha}(\tau)<N_{\beta}(\tau)$ or if, for a fixed $n$, $\text{perf}_{\alpha}(n)>\text{perf}_{\beta}(n)$.
3 Related Work
Distillation.
[…]
Synthetic question and answer generation.
[…]
Selective question and answer generation.
[…]
3.1 Assigning a Value to Data
Active learning.
Our work makes use of ideas from active learning, which seeks to maximise data efficiency by iteratively identifying and prioritising informative samples for labelling (Settles, 2009; Settles and Craven, 2008). Classic strategies for active learning include model prediction disagreement (Freund et al., 1997; Houlsby et al., 2011), uncertainty (MacKay, 1992; Gal et al., 2017; Kirsch et al., 2019), and dataset summarization (Sener and Savarese, 2018; Mirzasoleiman et al., 2020; Coleman et al., 2019). Effective, contemporary methods trade-off between predictive uncertainty and sample diversity in a fashion that is commensurate with large neural networks (Ash et al., 2021; Saran et al., 2023). We consider language model-aligned variations of two popular methods for active learning: uncertainty sampling (Settles and Craven, 2008), and BADGE, a more modern algorithm that trades-off between predictive uncertainty and the diversity of selected data (Ash et al., 2019).
Data selection.
Related methods aim to estimate the value of data to guide selection, typically using a labelled dataset $(x,y)$. Data can be valued using Shapley values (Ghorbani and Zou, 2019), influence functions Koh and Liang (2017) or by matching training data to evaluation datasets Just et al. (2023); Kessler et al. (2025); these methods have shown limited effectiveness for language modelling. LLMs have been used to score data points (Zheng et al., 2023) and for selecting question-answer samples for SFT (Liu et al., 2024b; Jiang et al., 2023b; Chen et al., 2023). Still, it has been shown that LLM scores exhibit biases that hinder their effectiveness in this setting (Xiong et al., 2024; Dorner et al., 2025; Panickssery et al., 2024). Alternative approaches use training loss or gradient norms with respect to student parameters as an estimate of learning progress (Loshchilov and Hutter, 2015; Katharopoulos and Fleuret, 2018; Jiang et al., 2019; Li et al., 2024; Mindermann et al., 2022; Evans et al., 2024; Dai et al., 2025). However, this has shown limited data efficiency for language models (Kaddour et al., 2023). Reward models are commonly used to select data points for SFT (Cao et al., 2023; Dubey et al., 2024). This work focuses on reward selection because of its popularity.
4 Iterative Synthetic Data Generation
The general iterative synthetic data generation process studied in this paper is shown in Algorithm 1 (Jiang et al., 2023b; Lee et al., 2024). We expand upon the algorithm’s design choices in the next sections. Most of these methods can be thought of as explicitly scoring each sample with a value $\{s_{i}\}_{i=1}^{n}$ where $n=|D_{0}|$ and $D_{0}$ is the initial question-answer seed dataset. In these cases, we can select $m=|\bar{D}_{t}|$ points with the highest scores equivalent to selecting the “hardest” points, with the highest uncertainty for instance (described in the next section), which is sometimes called “argmax” selection $\bar{D}_{t}=\text{argmax}_{m}\,\left\{s_{i}\right\}_{i=1}^{n}$. For completeness, we ablate these decisions, for example instead selecting the “easiest” points with lowest uncertainty, and sampling proportionally to scores instead of using argmax selection (Section C.6). Concrete instantiations of selection strategies $\phi$ are outlined below.
Algorithm 1 Iterative synthetic data generation algorithm for question and answer datasets.
Input: Seed dataset $D_{0}$, test set $D_{\text{test}}$, train set $\hat{D}_{-1}=\{\,\}$, student $f_{{\bm{\theta}}}(\cdot)$, selection algorithm $\phi$.
1:for $t=0,\ldots,T$ do
2:  Generate SLM predictions on $D_{t}$: $\{z_{i}=(x_{i},\hat{y}_{i})\}_{i=1}^{n}$ where $x_{i}\in D_{0}$ and $\hat{y}=f_{{\bm{\theta}}}(x)$.
3:  Select data subset: $\bar{D}_{t}=\phi(D_{t})$. $\triangleright$ See Section 4.1 for details.
4:  Generate synthetic dataset: $\hat{D}_{t}=\text{Generate}(\bar{D}_{t})$. $\triangleright$ See Section 4.2 for details.
5:  SFT on $f_{{\bm{\theta}}}(\cdot)$ using $\hat{D}_{t}:=\hat{D}_{t}\cup\hat{D}_{t-1}$ and evaluation on $D_{\text{test}}$.
6:end for
4.1 Selection Algorithms
Uncertainty sampling.
A common method in the active learning literature is uncertainty sampling, which, for non-sequential classification models, prioritizes data whose probability mass on the most likely class predicted by the model is smallest (MacKay, 1992). In the sequential, Transformer-based setting, we can score a data point with the loss of the response tokens under the student $f_{{\bm{\theta}}}(\cdot)$ with parameters $\bm{\theta}$} as $\mathcal{L}(z_{i},{\bm{\theta}})$ (Settles and Craven, 2008). When the targets used to produce a loss are the model’s own generations, this score reflects an uncertainty in the produced sequence. Note that our setting gives us access to the ground-truth label associated with $x$ as well, and thus allows us to compute a true loss in a fashion commensurate with conventional model training (Loshchilov and Hutter, 2015). Interestingly, we find this to be less effective empirically than using the former, uncertainty-based approach (Section C.6).
Reward scores.
[…]
LLM-as-a-judge scores.
We can also leverage the reasoning ability of an LLM teacher model to score an SLM’s predictions. This strategy asks the LLM teacher to score the detail, quality and correctness of the student’s answer and reasoning with a value between $[1,10]$. In particular, we use pairwise LLM-as-a-judge scoring which has been shown to be most effective (Zheng et al., 2023). Two separate answers are given for the teacher to decide which it prefers by providing scores for both: $s_{i}^{t},s_{i}=\texttt{LLM}(\hat{y}^{t}_{i},\hat{y}_{i},x_{i})$ where $\hat{y}^{t}_{i}=\texttt{LLM}(x_{i})$ is teacher’s answer, $s_{i}^{t}$ is the score for the teachers answer and $\hat{y}_{i}$ the student answer. This is expensive, as it requires the teacher to produce an answer in addition to scoring.
BADGE.
Batch Active learning by Diverse Gradient Embeddings (BADGE) is a two-stage active learning algorithm. It first represents all candidate data using the last-layer gradient of the loss induced by treating the generated sequence as ground truth, $\nabla_{{\bm{\theta}}_{o}}\mathcal{L}(\hat{y}=f_{{\bm{\theta}}}(x))$, where ${\bm{\theta}}_{o}$ are output-head parameters. BADGE then approximately samples from a $k$-DPP to identify gradients that are both high-magnitude and diverse (note that high-magnitude gradients are high-loss generations, suggesting high predictive uncertainty) (Ash et al., 2019). Like in uncertainty sampling, our setting allows us to use ground-truth target sequences, which would make these gradient representations of the sort used for optimization, but we found that using generated sequences resulted in better performance. Because the un-embedding layer of a Transformer is typically extremely large, we use a sparse random projection to efficiently reduce dimensionality while preserving geometric relationships (Johnson et al., 1984).
4.2 Prompt-based Synthetic Data Generation
Selected data points $\bar{x}_{i}\in\bar{D}_{t}$ are added to a synthetic data generation prompt for the LLM teacher model to generate a synthetic question $\hat{x}_{i}$ (Xu et al., 2024; Mitra et al., 2024; Jiang et al., 2023b; Lee et al., 2024). The teacher is then prompted to produce chain-of-thought reasoning and a final answer for $\hat{y}_{i}$. We generate a synthetic data point $\hat{z}_{i}=(\hat{x}_{i},\hat{y}_{i})$ using $\hat{x}=\texttt{LLM}(\bar{x}_{i})$ and $\hat{y}_{i}=\texttt{LLM}(\hat{x}_{i})$. So $\hat{D}_{t}=\text{Generate}(\bar{D}_{t})=\left\{\hat{x}_{i}=\texttt{LLM}(\bar{x}_{i}),\hat{y}_{i}=\texttt{LLM}(\hat{x}_{i})\right\}_{i=1}^{m}\text{, where}\,\bar{x}_{i}\sim\bar{D}_{t}$. For details on prompts used for each dataset see Section D.2.
5 Experiments
This section empirically probes the data efficiency of iterative synthetic data generation against static data generation, and provides recommendations for scoring and selection design choices for data efficiency. We find that prioritizing challenging data, as measured by the student’s loss on its own generations, to be at least as data efficient as teacher-based LLM scoring methods, and often more efficient.
LLM-based scoring can behave erratically, particularly for unusual tasks likely outside of the model’s training data distribution. Paired with the additional expense of using a large LLM to score data, more general approaches, like uncertainty sampling, appear to be more reliable and effective.
We further explore this improved data efficiency and show that on average synthetic data inherits some properties from samples used to generate them. If we select data that are difficult for the student—measured by a high loss or a low reward for example—the resulting synthetic data from the teacher is difficult as well, resulting in lower student accuracies on these generated samples than random selection.
At each iteration $t$ we use a given acquisition algorithm to select 11k samples $\bar{D}_{t}$ from $D_{t}$, before sending each to the teacher model to generate corresponding synthetic data $\hat{D_{t}}$. These data are appended to synthetic data from all previous iterations before reinitializing the student model and refitting its parameters with gradient descent.
5.1 Datasets
This section presents results on four distinct reasoning datasets in conjunction with four different models. GSM8k is a popular mathematics dataset comprised of school level maths problems (Cobbe et al., 2021), which we use with a Mistral-7B-Instruct-v0.3 student (Jiang et al., 2023a). Similarly, we include the more challenging Math1-3 dataset (Hendrycks et al., 2021), which is comprised of $5$ distinct levels of question difficulty—we use the easiest levels, $1$ to $3$, to finetune a Llama-3-8B-Instruct student (Dubey et al., 2024). We further experiment with the logical reasoning dataset ProntoQA (Saparov and He, 2023), composed of synthetically generated chain-of-thought style reasoning questions, with a Qwen1.5-7B-Chat student. Finally, we consider the Game of 24 dataset, which requires finding arithmetic operations to obtain $24$ given $4$ input integers. Here we use a Qwen2.5-7B-Instruct student (Qwen et al., 2025). Specifics are provided in Section D.1.
For all datasets except for Game of 24 we use prompt-based synthetic data generation with a GPT-4o teacher (prompts in  Section D.2). Instead, we use backward reasoning: if the answer is 13*8-10*8=24, for example, we can construct a new question by setting two integers to variables a*b-10*8=24 and solving to generate new questions (Jiang et al., 2024). We use a GPT-o3-mini teacher for backward reasoning, it qualitatively produces better question-response pairs than GPT-4o (Section D.2.4).
5.2 Finetuning Setup
To enable new instruction-following capabilities we finetune our student on synthetic data $\hat{D}_{t}$, which are appended to synthetic data from all previous iterations $\hat{D}_{<t}$. For efficient training we adapt LoRA layers (Hu et al., 2022) after each iteration of acquiring data and fitting the model. We avoid warm starting SFT parameters from their pre-trained values and instead use a fresh, random reinitialization (Ash and Adams, 2020; Springer et al., 2025). We set the LoRA rank and alpha parameters to the same value (see Section B.1) and adapt all linear layers. For optimization we use Adam (Kingma and Ba, 2014), clamp the gradient norm to a maximum of $2.0$, and use a batch size of $24$ with $2$ gradient accumulation steps. The learning rate decays linearly with a warm up period of $15\%$ of the total number of epochs. For Game of 24 we use a cosine decay learning rate schedule down to a minimum of $1\text{e-}9$ (Ni et al., 2025). During optimization we perform checkpointing and load the checkpoint with the best performance on a held-out validation set after SFT. We search over learning rates, LoRA ranks and the number of training epochs on this held-out validation set as well (Appendix B). We use a single 80Gb A100 or H100 GPU for all experiments.
Figure 2: SFT performance on 11k data points for various datasets and SLMs. We compare the effect of synthetic answer generation and synthetic question and answer generation to using the seed dataset, $D_{0}$ for SFT. 0-shot SLM and teacher performances are included for reference. All datasets use a GPT-4o teacher, for Game of 24 we use a GPT-o3-mini teacher. Using synthetic data, either as answers paired with real questions (syn. a) or both questions and answers (syn. q&a) improve performance past using the seed dataset alone (FT).
5.3 Algorithms
This paper considers a variety of selection algorithms. Prior work has shown that prioritizing “hard” samples accelerates learning (Section 3.1), which we also find to be the case for iterative synthetic data generation (Section C.6). This approach prioritizes high-uncertainty data, measured as the model’s loss on greedily decoded student generations, which we denote as “loss (high)” throughout this section. We also consider a low-reward selection algorithm (“rwd (low)”), also using the student’s own predictions, which scores generations using an external model. We use a Skywork-Reward-Llama-3.1-8B-v0.2 reward model which is the highest scoring 88b model on RewardBench (Lambert et al., 2025) at the time of writing.
We use Lion (Jiang et al., 2023b) as a baseline, which compares the student and teacher answer LLM-as-a-judge scores to classify each data point as either an easy or a hard before sampling equally from both sets. For completeness, we also consider a baseline that only samples from the hard set, denoted as LLM-as-a-judge (hard) (Jazbec et al., 2024). We use the same prompts for LLM-as-a-judge scoring as Jiang et al. (2023b).
We further consider prioritizing data with incorrect student answers, $s_{i}=\bm{1}\{\hat{y}_{i}\neq y\}$ as a proxy for prioritizing hard samples (Lee et al., 2024). In a similar spirit to Lion, we can instead sample evenly from correct and incorrect pools to maintain diversity in the seed data (Liu et al., 2024a). Since correct and incorrect scoring requires a verifier and ground-truth answers, we do not compare them to other scoring methods that do not use label information and instead place these supplementary results in Section C.1.
Figure 3: Student performance over successive synthetic data iterations with growing training sets. In all cases, selection based on uncertainty (loss) performs approximately as well as LLM-based scoring strategies (rwd and llm-judge), without requiring additional queries to an LLM. Further, for tasks that are out-of-distribution for the scoring model, like Game of 24, these mechanisms can perform even worse than random sampling. Horizontal lines in each inset plot denotes the proportion of data random sampling would require to achieve the same performance as the best active selection strategy in the corresponding experiment.
5.4 Results
This subsection presents our main results, which includes (1) performance comparisons between fitting the SLM on seed data and synthetic data (Section 5.4.1), and (2) between standard active learning strategies and more modern, LLM-based alternatives (Section 5.4.2). Further, we (3) analyse synthetically generated data to demonstrate it retains important properties of the original seed data, providing the underlying property of this mechanism affording the effectiveness of active methods (Section 5.4.3). Throughout this section, note that static generation is equivalent to random sampling of prompts for synthetic data generation in our setting, as it is not conditioned on the current state of the student. Unless stated otherwise, results show the mean and standard deviation over $3$ independent runs***Website and code: https://iterative-sd.github.io/.
5.4.1 Synthetic Data Improves Performance
SFT on synthetic data results in significantly improved capabilities when compared to using the original seed dataset.  Figure 2 compares SFT performance on the seed data to synthetic data of equal size, showing a dramatic increase in performance across all datasets when doing SFT on synthetic question-answers pairs. In the same figure, we see large increases in performance when using synthetic answers $z_{i}=(x_{i},\hat{y}_{i})$ instead of seed answers $y$, likely due to better formatting and high quality chain-of-thought in synthetic answers. In Game of 24 there is a small drop in
Figure 4: Pairwise winrate over all datasets and methods. ${\bm{P}}_{ij}$ corresponds to the number of times algorithm $i$ outperforms $j$. Overall performance is shown in the last row (lower is better).
performance when training on synthetic questions and answers compared to synthetic answers only, showing that the generation of novel questions by the teacher yields some lower quality synthetic questions. Regardless, next we show how this enables us to scale dataset sizes efficiently.
5.4.2 Iterative Generation is More Data Efficient than Static Generation
Active selection is more data efficient than random sampling for generating productive synthetic data, resulting in better performance using fewer samples. Figure 3 shows learning curves, with each plot measuring the test accuracy of a given selection method as a function of the labelling budget; each point is an active learning iteration. As mentioned earlier, random selection, because it is not conditioned on the current state of the student, is equivalent to the typical approach of static generation at the indicated data size. We find that this technique is often outperformed by a student-in-the-loop alternative—horizontal lines indicate the number of additional samples that would be required by static, random sampling in order to achieve the same performance as the best active learner in the corresponding plot (between $1.3\times$ and $2\times$). We find that uncertainty sampling performs roughly as well, and often better than LLM-based scoring methods. For some datasets, like Game of 24, reward-based methods do quite poorly, likely as a result of the task being out of distribution for the reward model.
Figure 4 aggregates performance differences between all selection strategies and model-dataset pairs considered in this paper. Each experiment composing this figure produced a learning curve in Figure 3, with each method producing a different SLM test accuracy for a variety of generation budgets. Here, we aggregate results by measuring which algorithms outperform their peers at each generation budget across all models and datasets.
Specifically, we aggregate results as a pairwise winrate matrix $\bm{P}$}. We increment ${\bm{P}}_{ij}$ if $\bm{1}\{\hat{\mu}_{i}-\alpha\cdot\hat{\text{se}}_{i}>\hat{\mu}_{j}+\alpha\cdot\hat{\text{se}}_{j}\}$, where $\hat{\mu}_{i}$ is the sample mean and $\hat{\text{se}}_{i}$ is the standard error of the performance of algorithm $i$ for a dataset, for a particular dataset size, and $\alpha$ is the confidence level which we set to $1$ (making it a $68\%$ confidence interval). By summing the “wins” across the rows and normalizing we can understand how often algorithms are outperformed on average. Column-wise averages are shown in the last row, where lower is better, to understand which algorithm is more data efficient in total. We find that random sampling—equivalent to static generation of data—is outperformed by various other methods that use the student model to guide synthetic data generation (Figure 4).
We can glean from Figure 4 that the highest-performing approach is simply uncertainty sampling, using the SLM’s loss on its own generations. LLM-as-a-judge also tends to be somewhat effective, though by a reduced margin. Interestingly, BADGE and Lion, which both aim to select diverse data, do not perform much better than random sampling (Figure 4). This is likely because synthetic data generation is noisy (Section 5.4.3), which diverse selection may exacerbate.
Because of the need to access a teacher model for scoring, LLM-as-a-judge is computationally demanding. Assuming that the cost of evaluating the teacher model dominates the cost of evaluating the student, a common assumption in the active learning literature and a reasonable assumption as the number of parameters of the teacher model can be $3$ orders of magnitude larger than the student models we consider. Then, if we consider the total number of teacher input and output tokens as a budget instead of the number of generated samples, Lion and LLM-as-a-judge (hard) are
Figure 5: Active learning curves on GSM8k: student performance against the number of teacher input and output tokens. The total number of input and output tokens are a proxy for the amount of compute used by the teacher for various methods.
far more expensive than other methods (Figure 5). Our results suggest this additional compute is better allocated towards simply generating more synthetic data with a cheaper and more effective selection strategy, like uncertainty sampling.
Figure 6: The rank correlations between original and synthetic dataset scores from iterative synthetic data generation. We plot student loss and reward scores and show Spearman’s rank correlations ($\rho$) between dataset medians before and after synthetic data generation. We zoom in on relationships at an individual data-point level where there is low correlation between the original and synthetic data point scores (centre). The red line is the line of best fit to the data. All rank correlations are highly significant ($p<0.001$).
Reward scoring also requires an external model, but because we can use a reward model that has the same number of parameters as our student, calls to the reward model are generally less expensive than to a teacher—we opt to not treat them in the same way and do not count the number of input tokens to the reward model in Figure 5. Overall random selection requires between $33\%$ to $100\%$ more SFT data to obtain the same performance as the best selection methods across all datasets (Figure 3). For $2/4$ of these datasets, iterative synthetic data generation using the loss on the student’s own predictions leads to more data efficient results compared to prior works that perform SFT using similarly sized datasets (see Section C.2 for comparisons).
5.4.3 Fidelity of Synthetic Data to its Original Data
Synthetic data generation is a noisy process, perturbing data by rephrasing, complicating or simplifying and adding chain-of-thought rationales. As such the score—either uncertainty (measured by the loss) or data quality (measured by the reward)—of an individual seed sample and its corresponding generated sample appear to have little to do with each other. In aggregate, however, we find high rank correlation between the median score of the seed and generated datasets (Figure 6). This relationship is the underlying principal governing why careful selection of the seed question-response pairs is important: Generated samples inherit underlying properties from the data used to produce them. These attributes, such as the SLM loss, shape the student’s optimization trajectory and generalization capabilities.
Figure 7: The percentage difference in synthetic data cumulative accuracy between samples ordered by score and randomly shuffled. Data are sorted either by uncertainty (high to low) or reward (low to high). Positive values suggest that score ordering picks more difficult synthetic samples in turn yielding lower accuracies. For each original data point we score it using the student model from the first and final iteration of iterative synthetic data generation (rows). See Figure 11 for the cumulative accuracies for individual replicates.
Unsurprisingly, samples with high uncertainty, again measured using the student’s loss on its own generations, tend to also be samples for which the student model struggles to obtain a correct answer. Figure 7 sorts samples by this uncertainty and compares the model’s accuracy on these data in comparison to a random shuffle, this is equivalent to random selection. Specifically, we plot the cumulative percent difference between the SLM accuracy on samples sorted by loss and the SLM accuracy on randomly ordered data. The curve in Figure 7 presents this as a function of the number of samples being used in the calculation, and is repeated for scores from a reward model. In both cases, prioritizing data in this fashion often effectively prioritizes low-accuracy samples, as indicated by the curve’s positive values. The trend is least clear for the ProntoQA dataset, which shows reward scoring as positive and uncertainty scoring as neutral or negative—unlike for other datasets, reward scoring was indeed a slightly more effective selection strategy (fig. 3). Low reward selection also selects lower-accuracy samples compared to random for the Game of 24 dataset (Figure 7), but still performs more poorly in terms of SFT performance than random (Figure 3). This is because the reward scorer introduces biases; specifically, it prefers longer responses, an effect that has been observed in other work (Shen et al., 2023; Bu et al., 2025).
6 Conclusion and Discussion
This work shows that student-in-the-loop synthetic data generation yields more data-efficient SLM improvements than static one-shot generation, and that simple, active-learning–inspired criteria for selecting seed examples outperform more elaborate LLM-based judging. We demonstrate that synthetic data are partially steerable, with teacher outputs reflecting the properties of selected seeds. These results highlight the value—and underexploration—of principled, effective selection strategies for advancing SLM training. Limitations are overviewed in Appendix A.
7 Acknowledgements
We thank Guoqing Zheng for insightful discussions.
Appendix
Appendix A Limitations
Iterative synthetic data generation for finetuning.
We only consider SFT, we do not consider efficient synthetic data generation to accelerate training for RLHF, continual pre-training [Yang et al., 2025] or pre-training [Maini et al., 2025], for instance. These are promising directions of future work.
The limits of the teacher.
We assume that the teacher is able to generate high quality questions and answers. For GSM8k, Math1-3 and ProntoQA the teacher performance is high and so we assume $\hat{z}_{i}$ is correct. For Game of 24 we rely on backward reasoning (specific to arithmetic) and a verifier to assess the teacher’s synthetic data. We have yet to test the limits of prompt-based synthetic data generation in settings where teacher capabilities fall short.
Data generation is noisy.
We can obtain improved student capabilities using iterative synthetic data generation. However, synthetic data generation is a noisy process where we show that properties of the selected datasets are preserved (Section 5.4.3). However, it is not clear how we can guarantee that synthetic data retains desirable properties from the seed dataset. For example, reward scoring performs poorly for the Game of 24 since it is biased by long student answers despite also selecting low quality student responses for synthetic data generation as required.
Model Dataset LoRA Rank Learning Rate Epochs Mistral-7B-Instruct-v0.3 GSM8k  seed 32 1e-4 10 Llama-3-8B-Instruct Math1-3  seed 32 1e-6 13 Qwen1.5-7B-Chat ProntoQA  seed 32 1e-5 13 Qwen2.5-7B-Instruct Game of 24  seed 16 1e-5 13 Mistral-7B-Instruct-v0.3 GSM8k  synthetic 32 1e-4 10 Llama-3-8B-Instruct Math1-3  synthetic 64 1e-4 13 Qwen1.5-7B-Chat ProntoQA  synthetic 32 1e-5 13 Qwen2.5-7B-Instruct Game of 24  synthetic 16 5e-4 30
Table 1: Optimal hyper-parameters for LoRA fine-tuning for all seed and synthetic datasets.
We have presented an initial study of the “steerability” of synthetic data generation. However the ability to add further desirable properties is left for future work.
Appendix B Additional Experimental Setup Details
We introduce additional details of our experimental setup from Section 5.3. We outline the hyperparameter grid search for SFT.
B.1 LoRA Hyper-parameter Tuning Setup
To obtain the best hyperparameters for our seed datasets $D_{0}$ and our synthetic datasets $\hat{D}_{t}$, we sweep through a grid of learning rates, number of training epochs and LoRA rank hyper-parameters using 11k question-answer pairs from the original seed dataset and 11k question-answer pairs synthetically generated by the teacher model. Refer to Table 1 for the optimal hyperparameters found in our sweep.
Appendix C Additional Results
We introduce additional results that support the claims in our main paper. In Section C.1, we introduce the results of prioritizing synthetic data generation using incorrect student predictions [Lee et al., 2024] and an even number of correct and incorrect student data [Liu et al., 2024a]. We do not include these results in the main paper for comparison since they require the ground truth answer $y$ for scoring unlike the other active scoring methods considered (Section 5.4.2). In Section C.2, we compare iterative synthetic data generation with comparable SFT methods from the literature. In Section C.3, we analyse the workings of synthetic data generation to show that despite introducing noise, the synthetic data retains the scores of the original selected seed data in aggregate. Furthermore, in Section C.4, we study how the synthetic datasets which are prioritized by low reward and high loss selection algorithms result in more difficult synthetic datasets since we observe lower student accuracies. In Section C.5, we show how the different scorers produce synthetic datasets with different token frequency distributions. These observations explain why selecting data prior to synthetic data generation results in data that has similar properties to our selected data and therefore enhanced student performance upon finetuning. Finally, in Section C.6, we compare various design choices for iterative synthetic data generation (Algorithm 1).
Figure 8: Iterative synthetic data generation learning curves, showing student SFT performance after training on synthetic data of increasing size with incorrect [Lee et al., 2024] and EvoKD [Liu et al., 2024a] data prioritization. Each consecutive increase in dataset size corresponds to an iteration of iterative synthetic data generation (Algorithm 1). Learning curves are across various dataset student model pairs. Curves are an average and standard error over 33 replicates.
C.1 Prioritizing Incorrect Samples
Figure 9: The pairwise win rate matrix over all datasets and all methods including incorrect prioritization. Element ${\bm{P}}_{ij}$ corresponds roughly to the number of times algorithm $i$ outperforms algorithm $j$ including results of incorrect student answer prioritization [Lee et al., 2024] and EvoKD [Liu et al., 2024a]. Column-wise averages at the bottom display overall performance (lower is better).
Prioritizing incorrect student predictions yields strong performance on all but one of the datasets we consider. A simple data point scoring mechanism is to assign a $\{0,1\}$ score for an incorrect or correct answer from the student model. This scoring mechanism requires a verifier or the ground truth answer $y$ and so is not directly comparable to the active scoring methods we consider that do not require the ground truth answer for scoring (Section 5.3). Regardless, we show the results of performing iterative synthetic data generation by prioritizing incorrect samples in Figure 8. For GSM8k this method severely underperforms other prioritization methods and random sampling.
Dataset Method LLM SFT Dataset Size Performance GSM8k Teacher GPT-4o n/a $94.9\pm 1.1$ Orca-Math [Mitra et al., 2024] Mistral-7B-Instruct-v0.3 1010k $70.2$ OpenMathInstruct [Toshniwal et al., 2024] Mistral-7B-Instruct-v0.3 1.81.8M $80.2$ Iterative Synthetic Data Generation (ours) Mistral-7B-Instruct-v0.3 1010k $80.6\pm 1.2$ Math1-3 Teacher GPT-4o n/a $91.8\pm 0.7$ Iterative Synthetic Data Generation (ours) Llama-3-8B-Instruct 1010k $56.1\pm 0.9$ ProntoQA Teacher GPT-4o n/a $98.9\pm 0.4$ Iterative Synthetic Data Generation (ours) Qwen1.5-7B-Chat 88k $96.9\pm 0.8$ Game of 24 Teacher GPT-o3-mini n/a $22.6\pm 1.8$ UFT [Ni et al., 2025] Qwen2.5-7B-Instruct 13.713.7k $30.2\pm 2.1$ Iterative Synthetic Data Generation (ours) Qwen2.5-7B-Instruct 66k $85.0\pm 1.3$
Table 2: Iterative synthetic data generation performs comparably to state-of-the-art SFT methods on certain datasets. The results of iterative synthetic data generation using high loss selection, as this selection method performs the best overall. We compare only to methods that use the same LLM and omit work that relies on larger datasets to achieve higher performance, as we cannot determine whether such gains stem from better techniques or simply from increased data. All SFT methods report the amount of data used for SFT. We report a mean and standard error over multiple seeds for our work, however some baselines only report a single seed.
For Math1-3 and Game of 24 incorrect student answer prioritization is as data efficient as high loss scoring which is the most data efficient method identified in Section 5.4.2. For the ProntoQA dataset incorrect answer prioritization obtains results on par with the best scoring methods if not better results for certain dataset sizes $n$. Considering a pairwise win-rate (described in Section 5.4.2) we can see from the row for incorrect prioritization that it is more data efficient in many instances with a high number of “wins” versus other methods. However at the same time looking at the corresponding column it is outperformed by many of the other methods in particular high loss and low reward selection due to its poor performance on the GSM8k dataset so it results in a poor overall score in the final row (Figure 9). Overall it is a simple method and has the possibility of obtaining strong capabilities and being more data efficient than random sampling in certain domains.
EvoKD underperforms random sampling and other active selection methods. Similar to Lion, which samples evenly from easy and hard data pools as determined by LLM-as-a-judge scores, we can sample data evenly from correct and incorrect student predictions for synthetic data generation [Liu et al., 2024a]. Evolving Knowledge Distillation (EvoKD), denoted as “EvoKD” in Figure 8, can be viewed as a diversity-based sampling approach for synthetic data generation. It achieves performance comparable to incorrect-data prioritization on GSM8k and Game of 24, but underperforms it on Math1-3 and ProntoQA. For GSM8k, EvoKD shares the same pathologies as promoting incorrect samples, they both underperform random sampling. EvoKD also underperforms methods that explicitly promote difficult samples (Figure 9), since it promotes hard samples through incorrect prioritization while simultaneously including easy samples to preserve the original data distribution. Overall, diversity-based criteria underperform approaches that emphasize difficult samples across the methods and datasets we study.
C.2 Comparing to Other SFT Methods
Iterative synthetic data generation obtains comparable results to state-of-the-art SFT methods on certain datasets. Table 2 compares the results of iterative synthetic data generation with high-loss selection to prior works in SFT which use the same LLM and similar dataset sizes. In our definition of data efficiency (Section 2), we can only properly compare against baselines that use the same model and that perform SFT on datasets of the same size, or if a baseline has a lower performance on a larger dataset size. Then we can conclude whether our method or the baseline is more data efficient, as defined in Section 2. If a baseline has better performance with a larger dataset size, then it is not possible to say whether the baseline we are comparing against or our method is more data efficient without scaling to the same dataset sizes. Since we cannot disentangle the performance improvements due to data quality or to increased dataset sizes. For GSM8k our work is more data efficient when compared to Orca-Math Mitra et al. [2024]. Also for Game of 24 our method outperforms state-of-the-art SFT baselines that use a Qwen2.5-7B-Instruct LLM Ni et al. [2025]. For the Math1-3 and ProntoQA datasets we did not find a comparable SFT methods to compare data efficiency with.
C.3 Synthetic Data Generation Preserves Properties of the Selected Data
At the dataset level synthetic data generation preserves properties of the original seed dataset. We score the selected seed dataset and take a median over scores and compare to the median score over the resulting synthetic data. If we do this for all iterations, we observe a very high rank correlation between median scores in Figure 6. This indicates that the scores across the iterative synthetic data generation curriculum are similar before and after synthetic data generation.
Figure 10: The scores for individual datapoints before and after 1 step of synthetic data generation. We consider the loss and reward of the student’s predictions and look at the individual data points scores across all datasets. The Spearman correlation measures the rank correlation before and after synthetic data generation. The red line shows the line of best fit to these data. The number of asterisks denotes the rank correlation’s p-value: *** indicates $p<0.001$.
When we look at the scores over individual data points and consider the score of a selected data point and the corresponding score of the synthetically generated datapoint, then we find there is a low but significantly greater than 0 rank correlation between reward and loss scores for all datasets (Figure 10).
Figure 11: The synthetic data cumulative accuracies when using random sampling and score ordering: high to low loss and low to high reward. For each original data point we score it using the student model from the first and final iteration of iterative synthetic data generation (alternating rows). Then we generate a synthetic data point. We compare the cumulative accuracy over the synthetic data when ordering data randomly versus ordering according to the loss and reward scores. We plot individual replicates as individual lines.
These two observations are consistent: synthetic data generation is preserving distributional factors such as dataset uncertainty (as measured by the loss over student predictions) and dataset quality (as measured by the reward over student predictions). But the noise from prompt-based synthetic data generation means that there is a low but significant correlation between scores at an individual data point level.
C.4 Prioritizing Difficult Data Creates Difficult Synthetic Data
The teacher produces difficult synthetic data when hard samples are prioritized by the student. We score seed data according to its loss or reward and then generate corresponding synthetic data. We obtain the cumulative accuracy of the synthetic data ordered by the original data scores. A random ordering corresponds to random sampling, while ordering the cumulative accuracy according to a high to low loss or low to high reward corresponds to prioritizing “difficult” data as we do in iterative synthetic data generation. For random sampling the cumulative accuracy versus the amount of data seen so far follows a diagonal line (Figure 11).
We plot the cumulative accuracy curves for synthetic data ordered from high to low original data loss (loss (high) ordering) in the first two rows and by low to high original data reward (rwd (low) ordering) in the final two rows of Figure 11. For GSM8k and Math1-3 the cumulative accuracy curves for synthetic data ordered using high to low original data loss and low to high reward are below random sampling so prioritizing data according to these scores results synthetic data that the student gets lower accuracies versus random sampling. The synthetic data is “harder” using these active learning approaches and these “hardness” qualities are integrated in the synthetic data the teacher generates. This is also seen for the first iteration for the Game of 24 dataset for both scorers. In contrast, in the final iteration the student is able to get a high accuracy on the synthetic data and so it is difficult to see any difference between random ordering and prioritizing according to a high loss or low reward. This is also the case for the ProntoQA dataset, for the first iteration we see high student accuracies for the synthetic data making comparison versus random sampling difficult, despite the reward scorer obtaining better performance than random on the ProntoQA dataset (Figure 8).
To obtain the cumulative difference plots presented in the main body of this manuscript in (Figure 7), we simply take the vertical distances between corresponding random sampling cumulative accuracies and the scorer cumulative accuracies in Figure 11 and aggregate across all replicates to obtain means and standard errors.
Figure 12: The total variation distances between token distributions of our synthetic data and the original seed datasets. We observe differences in the token distributions over the course of iterative synthetic data generation across for data selection algorithms, indicating differences in the synthetic datasets arise due to the different selection algorithms used.
C.5 Different Selection Algorithms have their own Selection Biases
The different selection algorithms we consider manifest as differences in the synthetic dataset distributions. When we compare the synthetic datasets to the original seed datasets over the course iterative synthetic data generation, then differences between selection algorithms are evident by looking at the token distributions in Figure 12. In particular, we measure the difference between two token distributions using the total variation distance (TVD): $\text{TVD}(P_{D_{0}},P_{\hat{D}_{t}})=\frac{1}{2}\sum_{x\in V}|P_{D_{0}}(x)-P_{\hat{D}_{t}}(x)|$ where $x$ is a token in the vocabulary $V$ and $P$ is the empirical token distribution. The token distribution $P$ can be thought of as a histogram where the bin size is the normalized frequency of the token in the dataset. This distance is essentially looking at the absolute differences in token counts between two datasets. When measuring the TVD between synthetic datasets and the original seed dataset prior to selection, $D_{0}$. We can see that the distance varies between different selection algorithms which shows that there are differences in the synthetic datasets at a token distribution level. The Game of 24 dataset is the sole case where the selection algorithms yield almost indistinguishable TVDs, as its questions and answers draw from a highly restricted token range to compute 24 from four numbers using basic arithmetic operations. This points to there being distributional differences between synthetic datasets of different selection algorithms and thus shows that the selection algorithms manifest in different synthetic datasets with different properties over the course of iterative synthetic data generation. These distributional differences lead to performance differences between different selection algorithms which have been studied in the main results (Figure 4).
C.6 On the Design Choices for Iterative Synthetic Data Generation
Argmax selection, rather than sampling, results in the best SFT performance. In Figure 13, we compare various data prioritization design choices. The performance for scorers that prioritize data where the student answer is the most uncertain (high loss) or worse quality (low reward) results in the best performance when compared to data for which the model is confident (low loss) or is of better quality (high reward). Furthermore, we compare whether using the ground truth answer $y$ (denoted “gt” in Figure 13) or the student’s own prediction $\hat{y}$ is more data efficient. We can see worse performance when computing scores with the ground-truth answer for the loss scorer, while scoring with the reward model results in equal SFT performance. There is no benefit to using the ground truth answers over the student’s own predictions.
Figure 13: Performance of iterative synthetic data generation on various data scoring and selection options. We train on 11k data points at each iteration with a Mistral-7B-Instruct-v0.3 student on GSM8k. We compare prioritizing “difficult” or “easy” data points with a high or low loss or reward. We compare using ground truth answers $y$ to the student’s own predictions $\hat{y}$ and using argmax selection against sampling.
Finally, we compare selection methods: argmax selection and sampling and can see lower SFT performance when using sampling (labelled with “sampling” in Figure 13). We sample $m$ points by sampling from a distribution proportional to these scores: $\bar{D}_{t}\overset{m}{\sim}{\text{softmax}}(\left\{s_{i}\right\}_{i=1}^{n})$. We found poor performance when sampling because sampling from the softmax distribution of loss or reward scores results in a similar distribution of scores for selected data as if we performed random sampling. Moreover, if we select m=1m=1k data points from the GSM8k seed dataset and look at the distribution of loss scores via sampling for the highest and lowest 11k scoring data, then the distributions are indistinguishable to the naked eye. Argmax selection however produces distinct distributions (Figure 14).
Appendix D Dataset Further Details
In this section we provide in depth details on the datasets used in our experiments together with the dataset sizes used throughout our empirical study of iterative synthetic data generation (Section D.1). Also we provide the teacher prompts used for synthetic data generation (Section D.2).
We introduce the seed question and answer datasets $D_{0}$. The validation and test sets are taken from the original seed datasets as opposed to using synthetic data. The train sets $\hat{D}_{t}$ are synthetically generated. We summarize the datasets sizes in Section D.1. Unless otherwise stated we use a GPT-4o teacher. We prompt the teacher with few-shot examples from $D_{0}$ to generate a new synthetic questions [Liu et al., 2024a]. For all datasets we throw away similar synthetic questions if the rouge-score [Lin, 2004] with respect to all previously generated questions is above $0.7$ [Jiang et al., 2023b].
Figure 14: Distribution of losses for different sampling methods. We select 11k according to a high or low loss sampling (left) and argmax selection (right) for GSM8k and can see almost no difference when using sampling.
GSM8k.
We perform SFT on a Mistral-7B-Instruct-v0.3 [Jiang et al., 2023a] student on school level mathematics questions [Cobbe et al., 2021]. We use an external language model gpt4o-mini to assess whether the student’s answer is equivalent to the ground truth answer, in a similar manner to Mitra et al. [2024], see Section D.3 for prompting details. We take $748$ question-answer pairs from the test set as a validation set and use $500$ question-answer pairs as a test sethttps://huggingface.co/datasets/openai/gsm8k.
Math1-3.
[…]
ProntoQA.
The questions are synthetically generated logical chain-of-thought style reasoning questions with boolean answers [Saparov and He, 2023]. We perform SFT with a Qwen1.5-7B-Chat student model. We use an external language model gpt4o-mini to assess whether the student’s reasoning steps are correct and whether the student answer is equivalent to the ground truth answer like for GSM8k, see Section D.3 for details. We use $300$ question-answer pairs as a validation set and the remaining $200$ as a test set§§§We usehttps://huggingface.co/datasets/renma/ProntoQA for validation and testing, as a train set we use https://huggingface.co/datasets/longface/prontoqa-train like in [Huang et al., 2024], questions are distinct between these two ProntoQA datasets..
Game of 24.
[…]
Dataset Seed Size Validation Size Test size
GSM8k 7473 748 500
Math1-3 3504 500 500
ProntoQA 2880 300 200
Game of 24 2217 500 300
Table 3: Summary of the seed dataset sizes, validation and test set sizes. For all datasets we use 11k data points per iteration for finetuning.
D.1 Seed Dataset Sizes
We summarize the seed dataset sizes for all datasets used in our experiments in Table 3. The seed dataset $D_{0}$, is used for scoring and selecting data points to get the selected data $\bar{D}_{t}$. The selected data is then put forward for prompt-based synthetic data generation (Section 4.2). We set the validation and test sets to be from the original seed datasets. We use the resulting synthetic datasets $\hat{D}_{t}$ for SFT, we generate a fixed sized synthetic dataset to enable fair comparison between selection methods and assess data efficiency (Section 5.1).
D.2 Synthetic Data Generation Prompts
We provide the prompts used for prompt-based synthetic data generation (described in Section 4.2) below for all datasets used in our experiments:
D.2.1 Grade School Maths
Below is the prompt we use for synthetic question generation for GSM8k using a GPT-4o teacher. In the prompt below $\{0\}$ are few-shot examples of questions and answers: $\{z_{i}\}_{i=1}^{k}\sim D_{0}$, we set $k=5$ for all our experiments and $\{1\}$ is the question from the data selected by the student: $\bar{x}=\bar{z}[0]$ where $\bar{z}\sim\bar{D}_{t}$. The few-shot examples are formatted as follows: #Given Instruction#: {} #Answer#: {}
GSM8k synthetic question generation prompt I want you to act as Instruction Creator.
Your objective is to rewrite a #Given Instruction# into a more complex version, to make it a bit harder.
The #Rewritten Instruction# must be reasonable and must be understood and responded to by humans.
Here are some #Examples#:
$\{0\}$
I want you to act as Instruction Creator.
Your objective is to rewrite a #Given Instruction# into a more complex version, to make it a bit harder.
The #Rewritten Instruction# must be reasonable and must be understood and responded to by humans.
You MUST complicate the #Given Instruction# using the following method:
1. Change the names of people #Given Instruction#.
2. Change the objects in the #Given Instruction#.
3. Change any quantities and durations in the #Given Instruction#.
4. Add 1 to 3 more operations in #Rewritten Instruction#.
5. Change the operations, for example: multiplication, division, subtraction, addition, percentages, fractions and combinations of these.
6. You should try your best not to make the #Rewritten Instruction# become verbose, #Rewritten Instruction# can only add 10 to 20 words into #Given Instruction#.
Use #Examples# to complicate #Given Instruction#.
’#Given Instruction#’, ’#Rewritten Instruction#’, ’given instruction’ and ’rewritten instruction’ are not allowed to appear in #Rewritten Instruction#.
#Given Instruction#:
$\{1\}$
#Rewritten Instruction#:
We use the following prompt to obtain synthetic answers from our GPT-4o teacher (and from our student model):
GSM8k answer prompt Question: {} Solve the problem step-by-step. Answer:
D.2.2 Math1-3
Below is the prompt we use for synthetic question generation for Math1-3 using a GPT-4o teacher, $\{0\}$ are few shot examples of questions, answers and the type of problem e.g. Geometry, Algebra etc. The number of few-shot examples is $5$ and are of the same type as the seed question. In the prompt below $\{1\}$ is the type of mathematics problem and $\{2\}$ is the question from the selected dataset: $\bar{x}=\bar{z}[0]$ where $\bar{z}\sim\bar{D}_{t}$. The few-shot examples are formatted as follows: The type of math problem is {}. #Given Instruction#: {} #Answer#: {}
Math1-3 synthetic question generation prompt I want you to act as an Instruction Creator for {1} mathematics problems.
Create a new question #Rewritten Instruction# by using #Given Instruction# as inspiration. The new question should have a single unique answer.
Ensure that the type of the question you generate #Rewritten Instruction# matches the type of instruction #Given Instruction#.
Make #Rewritten Instruction# different from #Given Instruction#.
The #Rewritten Instruction# must be reasonable, have a solution and must be understood and responded to by humans. Here are some #Examples#:
{0}
Use #Examples# as inspiration to make #Rewritten Instruction# different to #Given Instruction#.
’#Given Instruction#’, ’#Rewritten Instruction#’, ’given instruction’ and ’rewritten instruction’ are not allowed to appear in #Rewritten Instruction#.
#Given Instruction# is a {1} math problem.
#Given Instruction#:
{2}
#Rewritten Instruction#:
We use the following prompt for obtaining synthetic answers from our GPT-4o teacher (and for obtaining answers from our student model):
[…]
D.2.3 ProntoQA
We present the prompt we use for synthetic question generation using a GPT-4o teacher for the ProntoQA dataset [Saparov and He, 2023]. A datapoint from the ProntoQA dataset is comprised of a context, question and answer $z=(x=(c,q),y)$ where $x$ is comprised of the context $c$ and question $q$. The answers $y$ are boolean. The few-shot question generation is therefore comprised of contexts and questions for the teacher to generate new synthetic context and questions, $\hat{x}$. In the prompt below $\{0\}$ are few-shot examples of questions and answers from $\{z_{i}\}_{i=1}^{k}\sim D_{0}$, we set $k=5$ for all our experiments and $\{1\}$ is the question from the selected dataset $\bar{x}=\bar{z}[0]$ where $\bar{z}\sim\bar{D}_{t}$. The few-shot examples $\{0\}$ are formatted as follows: Context: {} Question: {}.
ProntoQA synthetic question generation prompt I want you to act as an Instruction Creator for logical problems.
Create a new question #Rewritten Instruction# by using #Given Instruction# as inspiration.
Make #Rewritten Instruction# different from #Given Instruction# by changing the names, objects and adjectives. Also vary the number of logical reasoning steps in #Rewritten Instruction#. Ensure that it is possible to answer the question with true or false answer.
The #Rewritten Instruction# must be reasonable, have a solution and must be understood and responded to by humans.
Here are some #Examples#:
{0}
Use #Examples# as inspiration to make #Rewritten Instruction# different to #Given Instruction#.
’#Given Instruction#’, ’#Rewritten Instruction#’, ’given instruction’ and ’rewritten instruction’ are not allowed to appear in #Rewritten Instruction#.
#Given Instruction#:
{1}
#Rewritten Instruction#:
We use the following prompt for obtaining synthetic answers from the GPT-4o teacher (and for obtaining answers from our student model):
ProntoQA answer prompt Context: {} Let’s think step by step. Response:
D.2.4 Game of 24
Below is the prompt we use for synthetic question generation using GPT-o3-mini teacher for the Game of 24 dataset. A datapoint from the Game of 24 dataset is comprised of a set of four numbers and the arithmetic one-line solution to obtain $24$. In the prompt below $\{0\}$ is the question, a set of numbers for instance $\bar{x}=[8,8,10,12]$ and $\{1\}$ is the arithmetic answer for instance $\bar{y}=(12-10)\times 8+8$ where $\bar{z}=(\bar{x},\bar{y})$ and $\bar{z}\sim\bar{D}_{t}$. We use backward reasoning to to obtain a new question and answer to the Game of 24 (see the prompt below). We verify that the synthetic answer evaluates to $24$ and that all the numbers from the synthetic question are also present in the synthetic answer. Since backward reasoning for synthetic data generation produces both the question and the answer, we then prompt our teacher, GPT-4o in a second step, with both the synthetic question and answer to get a synthetic reasoning trace without any verification of the reasoning steps to construct our synthetic dataset $\hat{D}_{t}$ (in the second prompt below).
Game of 24 synthetic question generation prompt I want you to act as an instruction creator. I want you to write a new problem to the game of 24.
The numbers {0} need to be used to obtain the number 24. Use each number once, even if a number is repeated use it multiple times, with the arithmetic operations +, -, *, / to obtain 24. Here is how the above numbers {0} are used to obtain 24: {1}.
I want you to create a new problem to the game of 24 using {1}. Let’s use a backward thinking method. Take two of the distinct numbers in {1}. Call them a and b. Then construct an equation with two unknowns, a and b. Pick integer values for the first variable b then solve for a.
For example the numbers 8, 8, 10, 13 can be used to get 24: 13*8-10*8=24. We can construct the following equation a*b-10*8=24 by substituting a=13 and b=8. Rearranging we get a=104/b. Let’s pick an integer which divides into 104 for b: b=4 therefore a=26.
[…]
Game of 24 prompt for synthetic reasoning steps Use numbers and basic arithmetic operations (+ - * /) to obtain 24. Each step, you are only allowed to choose two of the remaining numbers to obtain a new number.
Input: 4 4 6 8
Steps:
4 + 8 = 12 (left: 4 6 12)
6 - 4 = 2 (left: 2 12)
2 * 12 = 24 (left: 24)
Answer: (6 - 4) * (4 + 8) = 24
Input: 2 9 10 12
Steps:
12 * 2 = 24 (left: 9 10 24)
10 - 9 = 1 (left: 1 24)
24 * 1 = 24 (left: 24)
Answer: (12 * 2) * (10 - 9) = 24 Input: 4 9 10 13
Steps:
13 - 10 = 3 (left: 3 4 9)
9 - 3 = 6 (left: 4 6)
4 * 6 = 24 (left: 24)
Answer: 4 * (9 - (13 - 10)) = 24
Input: 1 4 8 8
Steps:
8 / 4 = 2 (left: 1 2 8)
1 + 2 = 3 (left: 3 8)
3 * 8 = 24 (left: 24)
Answer: (1 + 8 / 4) * 8 = 24
Input: 5 5 5 9
Steps:
5 + 5 = 10 (left: 5 9 10)
10 + 5 = 15 (left: 9 15)
15 + 9 = 24 (left: 24)
Answer: ((5 + 5) + 5) + 9 = 24
Input: {question}
Here is the final answer: {answer}
[…]
We use the following prompt to get answers from the student, similarly to Ni et al. [2025]:
Game of 24 student prediction prompt Use numbers and basic arithmetic operations (+ - * /) to obtain 24. Each step, you are only allowed to choose two of the remaining numbers to obtain a new number.
Input: 4 4 6 8
Steps:
4 + 8 = 12 (left: 4 6 12)
6 - 4 = 2 (left: 2 12)
2 * 12 = 24 (left: 24)
Answer: (6 - 4) * (4 + 8) = 24
Input: 2 9 10 12
Steps:
12 * 2 = 24 (left: 9 10 24)
10 - 9 = 1 (left: 1 24)
24 * 1 = 24 (left: 24)
Answer: (12 * 2) * (10 - 9) = 24 Input: 4 9 10 13
Steps:
13 - 10 = 3 (left: 3 4 9)
9 - 3 = 6 (left: 4 6)
4 * 6 = 24 (left: 24)
Answer: 4 * (9 - (13 - 10)) = 24
Input: 1 4 8 8
Steps:
8 / 4 = 2 (left: 1 2 8)
1 + 2 = 3 (left: 3 8)
3 * 8 = 24 (left: 24)
Answer: (1 + 8 / 4) * 8 = 24
Input: 5 5 5 9
Steps:
5 + 5 = 10 (left: 5 9 10)
10 + 5 = 15 (left: 9 15)
15 + 9 = 24 (left: 24)
Answer: ((5 + 5) + 5) + 9 = 24
Input: {question}
[…]
D.3 Evaluation prompts
To assess whether the student’s prediction is equal to the ground-truth answer we use gpt4o-mini to verify the correctness of the student. We use the following prompt and a system prompt which is different for each dataset used:
GSM8k and ProntoQA evaluation prompt Question:{} Problem Setter’s answer:{} Student’s answer:{}
For GSM8k we use the following system prompt for evaluation, similarly to Mitra et al. [2024]:
GSM8k evaluation system prompt As an expert Math teacher, your role is to evaluate a student’s answer to a word problem. The problem is accompanied by a correct solution provided by the problem setter. It is important to remember that there may be various methods to solve a word problem, so the student’s steps might not always align with those in the problem setter’s solution. However, the final answer, typically a number, should be unique and match the problem setter’s answer. Your task involves analyzing the student’s solution to identify any mistakes and determine whether the answer can be modified to correct the error. If the student’s answer is unfixable, consider creating practice problems to help improve their understanding. Use the following format: Error Analysis: In one sentence, extract the final answer from the problem setter’s solution and compare it with the student’s answer. Do they match? Final Verdict: Correct/Incorrect.
For ProntoQA we use the following system prompt for evaluation:
ProntoQA evaluation system prompt You are a logical expert. Your role is to evaluate a student’s answer to a logical reasoning problem. The problem is accompanied by a correct solution provided by the problem setter. Your task is to assess whether the problem setter’s answer and the student’s answer match. Use the following format: Error Analysis: In one sentence, extract the final answer from the problem setter’s solution and compare it with the student’s answer. Do they match? Final Verdict: Correct/Incorrect.
If the output contains string variations of "Final Verdict: Correct" then the student’s prediction is correct and wrong otherwise. For the Math1-3 and Game of 24 datasets we use pattern matching to extract the student’s answer and compare to the ground truth, see Section 5.1 for details.